8. Geo-M is considering adding a cash discount to its credit terms. If Geo-M offers 3/11 net 30 rather than its current net 30 policy, what annualized rate is the company charging customers who do not take the discount? Assume a 365-day year.

9. Fickle Sickles collects 15,000 checks per 365-day year with average amount $170 and total delay 5 days. A lockbox system would reduce that delay to 3 days, and it would also reduce FISI's check processing costs by $.01 per check. The annual fee on the lockbox would be $2,500. What is the annual net benefit associated with using the lockbox? FISI has an 7% cost of capital.
Answer #1
8] The 3% discount is the cost for the 19 days of credit.
Hence, cost of not taking discount for 19 days = 3/97 = 3.09%
Annualized cost [without compounding] = (3/97)*(365/19) = 59.41%
Effective interest rate [with compounding] = (1+3/97)^(365/19)-1 = 79.52%
9] Annual fee on the lockbox $      -2,500.00
Processing costs = 0.01*15000 = $         -150.00
Interest saved = (5-3)*(15000/365)*170*7% = $           978.08
Net annual benefit associated with the lockbox $      -1,671.92